OpenStudy (anonymous):

The cost for Jan to set up one internet connection and one cable connection in her house is $75. The cost for her to set up one internet connection and two cable connections is $100. How much does one cable connection cost?

OpenStudy (anonymous):

ok do you know how to start this?

OpenStudy (anonymous):

No idea.

OpenStudy (anonymous):

alright, well, we can start by writing a system of equations. call the internet connection "i" and the cable connection "c" what equations can you write with these two variables?

OpenStudy (anonymous):

i + c = 75

OpenStudy (anonymous):

alright, good, and there's one more - look at the second sentence

OpenStudy (anonymous):

i + c = 100? Idk

OpenStudy (anonymous):

close, but it's i + 2c = 100, because it's *two* cable connections. got it?

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

okay so have you learned about either the elimination method or the substitution method?

OpenStudy (anonymous):

okay, which one is easier for you/ makes more sense to you?

OpenStudy (anonymous):

Doesn't matter to me.

OpenStudy (anonymous):

ok then, solve for i in the first equation

OpenStudy (anonymous):

i = c - 75

OpenStudy (anonymous):

*i = 75 - c like this:  makes sense?